4. A block of mass 0.25 kg is placed on a vertical spring of constant k- 5000 N/m, and is pushed downward compressing the spring a distance of 0.1 m. As the block is released, it leaves the spring and continues to travel upward. To what maximum height above the point of release does the block rise?
Answer #1

mass = m = 0.25 kg

spring constant = k = 5000 N/m

compression = x = 0.1 m

from conservation of energy    1/2 k x^2 = m * g * h

                                                     0.5 * 5000* 0.1^2 = 0.25 * 9.8 * h

                                                               25 = 2.45 h

               height = h = 25 / 2.45 = 10.2 m